Solve the system of inequalities graphically:
y is greater than or equal to x-4
y is less than or equal to -2x-8

Thank you for your help!!

Answer by khai(18)

y is greater than or equal to x-4 -------------- equation (1)
y is less than or equal to -2x-8 -------------- equation (2)
multiply by 2 to equation (1)
2y is greater than or equal to 2x-8 ------------ equation (3)
add equation (3) and (2)
2y+y = (2x-8)+(-2x-8)
3y = -16
y = -16/3
substitute y value into equation (1)
x = -4/3